OpenStudy (anonymous):

determine the quadratic function f whose graph is given, the vertex is (2,-7) and y intercept is -3. quick can someone help me fast with this. Im being timed. Thanks

OpenStudy (anonymous):

what is the formula for a parabola

terenzreignz (terenzreignz):

Parabola is given by a quadratic function, like f(x) = ax² + bx + c Its vertex (h, k) is given by (-b/2a, f(-b/2a))

OpenStudy (anonymous):

so what will it be when the vertex is (2,-7) and y intercept of -3

terenzreignz (terenzreignz):

\[-\frac{b}{2a}=2\]\[\frac{4ac-b^{2}}{4a}=-7\] Could you work these out?

terenzreignz (terenzreignz):

By the way, y-intercept is just f(0)

OpenStudy (anonymous):

dont know, do not have time right now. I need answer to my problem right now. can you just write it out for me so i can understand better, please

terenzreignz (terenzreignz):

The y-intercept f(0) = -3, This can only mean f(0) = a(0)² + b(0) + c = -3 Meaning c = -3 Now, just find a and b.

OpenStudy (anonymous):

is it a=2 and b=-7

terenzreignz (terenzreignz):

If that were so, then the x-coordinate of your vertex (h) would be equal to -(-7)/2(2) = 7/4 = 1.75 Which is not equal to 2, as it should be, as it has been given. Use systems of equations.

OpenStudy (anonymous):

ok so how do i find a an b then?

OpenStudy (anonymous):

its been a long time for me,sorry

terenzreignz (terenzreignz):

Just use systems of equations.